OpenStudy (anonymous):

Solve the system of equations below x-2y=3 2x-3y=0 (3,9) (9,3) (-3,9) (-9,3)

OpenStudy (anonymous):

Multiply the top equation by -2

OpenStudy (anonymous):

It would be D then right?

OpenStudy (yttrium):

I think it is not.

OpenStudy (anonymous):

How would I solvje the equation?

OpenStudy (anonymous):

i have (-9, -6)

OpenStudy (yttrium):

We have x-2y = 3 and 2x-3y = 0 Therefore, using the second equation we can say that 2x = 3y and therefore, x = 3y/2 Substituting the x we got from the second eqn to the 1st eqn we'll get \[\frac{ 3y }{ 2 } - 2y = 3\] Upon simplifying 3y-4y = 6 -y = 6 y = -6 x = -9

OpenStudy (yttrium):

we both arrive at the same answer with @whalexnuker

OpenStudy (anonymous):

There is no -9 -6 option

OpenStudy (anonymous):

That much is apparent, lol

OpenStudy (yttrium):

That means, none among the choices is correct.

OpenStudy (anonymous):

x-2y=3 2x-3y=9 That's the right question, sorry. It would be D right?

OpenStudy (anonymous):

(9, 3) @mikeshaw

OpenStudy (anonymous):

So B

OpenStudy (anonymous):

x - 2y = 3 -->(-2) 2x - 3y = 0 --------------- -2x + 4y = -6 (result of multiplying by -2) 2x - 3y = 0 ---------------add y = -6 now sub -6 in for y in either of the original equations x - 2y = 3 x - 2(-6) = 3 x + 12 = 3 x = 3 - 12 x = -9 answer is (-9,-6) hmmm... let me try it a different way.... x - 2y = 3 --> x = 2y + 3 now sub 2y + 3 in for x in the other equation 2x - 3y = 0 2(2y + 3) - 3y = 0 4y + 6 - 3y = 0 y + 6 = 0 y = -6 now sub -6 in for y 2x - 3y = 0 2x - 3(-6) = 0 2x + 18 = 0 2x = -18 x = -9 same answer : (-9,-6)

OpenStudy (anonymous):

@kelliegirl33 We've all given him the same thing.

OpenStudy (anonymous):

oh wait...you changed the question... x - 2y = 3 -->(-2) 2x - 3y = 9 ---------------- -2x + 4y = -6 (result of multiplying by -2) 2x - 3y = 9 ----------------add y = 3 now sub 3 in for y 2x - 3y = 9 2x - 3(3) = 9 2x - 9 = 9 2x = 9 + 9 2x = 18 x = 9 answer (9.3)

OpenStudy (anonymous):

Its B